# Guarded Query Routing for Large Language Models

### v1

@@ 7 Discussion
Despite being designed explicitly for safety, existing guardrail techniques displayed poor performance in GQR-Bench, as they are not well-suited for ID routing.
Routing based on embedding similarity is an interesting direction, as it requires only few examples per domain. However, SemanticRouter’s GQR-Score was only moderate due to very limited OOD detection capabilities – which can be explained by the absence of any training examples for the different types of OOD queries.
@@
Our findings suggest that FastText and WideMLP can be attractive alternatives to LLM-based routers for production systems. They provide GQR-Scores of 80% with sub-millisecond latencies. Although LLM-based approaches (GPT-4o-mini and Llama3.1-8B) achieve higher GQR-Scores, they are 2–3 orders of magnitude slower than efficient alternatives.
Adopting these efficient text classifiers for query routing can help practitioners achieve substantial cost savings and latency improvements without compromising accuracy. This is particularly relevant for systems that handle large volumes of queries, where modest savings in computation per query can translate to immense benefits in terms of total cost and energy consumption.

### v3

@@ 7 Discussion
Despite being designed explicitly for safety, existing guardrail techniques displayed poor performance in GQR-Bench, as they are not well-suited for ID routing. Routing based on embedding similarity is an interesting direction, as it requires only a few examples per domain. However, SemanticRouter’s GQR-Score was only moderate due to very limited OOD detection capabilities – which can be explained by the absence of any training examples for the different types of OOD queries.
@@
Our findings suggest that FastText and WideMLP can be attractive alternatives to LLM-based routers for production systems. They provide GQR-Scores of 80% with sub-millisecond latencies. Although LLM-based approaches (GPT-4o-mini and Llama3.1-8B) achieve higher GQR-Scores, they are 2–3 orders of magnitude slower than efficient alternatives. Adopting these efficient text classifiers for query routing can help practitioners achieve substantial cost savings and latency improvements without compromising accuracy. This is particularly relevant for systems that handle large volumes of queries, where modest savings in computation per query can translate to immense benefits in terms of total cost and energy consumption.
@@ 8 Conclusions
Limitations and Future Work
The in-distribution datasets of GQR-Bench only differ in their domain (law, finance, health), where a variety of other aspects could be included, e.g., the complexity of a query (route it to a more powerful model), its language (route it to a multilingual model), or a particularly important query (route it to a human chat partner). All these aspects would be very interesting to study in future work, yet here we instead focused on variety in testing the robustness of the system with respect to out-of-distribution queries, covering other domains, other languages, and unsafe inputs.